In a common emitter transistor amplifier the load resistance is 5000ohm and input resistance is 1000ohm. If the peak value of input signal voltage is 10mv peak value of output voltage is (beta=100)
1
Expert's answer
2021-03-14T19:13:26-0400

Voltage gain can be found as follows:


"A_V=\beta\dfrac{R_L}{R_I}=\dfrac{V_O}{V_I},""A_V=100\cdot\dfrac{5000\ \Omega}{1000\ \Omega}=500."

Finally, we can find the peak value of output voltage:


"V_O=V_IA_V=10\cdot10^{-3}\ V\cdot500=5\ V."